OpenStudy (anonymous):

1/3 divided by 5/6 = ?

OpenStudy (anonymous):

2/5

OpenStudy (anonymous):

2/3 2/5 2/6 1/5

OpenStudy (alexandervonhumboldt2):

\[\frac{ 1 }{ 3 }/\frac{ 5 }{ 6 }=\frac{ 1 }{ 3 }*\frac{ 6 }{ 5 }=\frac{ 6 }{ 15 }=\frac{ 3*2 }{ 3*5 }=\frac{ 2 }{ 5 }\]